find two equations of lines that are parallel to the lines with these equations;
y=-7x+8
y=3/4x+3

Answer by ilana(307)

These lines are in slope-intercept form, y=mx+b, where m is the slope and b is the y-intercept. Parallel lines have the same slope but different y-intercepts (imagine that same line shifted up or down along the y-axis. So m should stay the same but b should change. For example, if the line is y=-7x+8, then some parallel lines are y=-7x+1, y=-7x-2, y=-7x+1112, y=-7x+34.34534534, y=-7x-343.543453. Enjoy:)